you invest $400. After one year, the total of the investment is $414.40. Use the formula A=P+Prt to find the annual simple interest rate for the investment, where A is the total of hte investment, P is the principal (amount invested), r is the annual simple interest rate, and t is the time in years.
Answer by checkley77(12844)

414.40=400+400*r*1
414.40=400+400*r
414.40-400=400t
14.40=400t
t=14.40/400
t=.036 or 3.6% is the interest rate.